The angular position of a swinging door is described by θ(t)= (5.00+ 2.00t²) rad




Determine the angular position, angular speed, and angular acceleration of the door at t=3.0s.


1
Expert's answer
2022-01-17T09:59:00-0500

"\theta(t)=5+2t^2,"

"\theta(3)=23,"

"\dot \theta(t)=4t,"

"\dot \theta(3)=12,"

"\ddot \theta(t)=4,"

"\ddot \theta(3)=4."